OpenStudy (anonymous):

A rectangular storage container with an open top is to have a volume of 10 m3. The length of this base is twice the width. Material for the base costs $15 per square meter. Material for the sides costs $9 per square meter. Find the cost of materials for the cheapest such container. (Round your answer to the nearest cent.)

jhonyy9 (jhonyy9):

what is your opinion ?

jhonyy9 (jhonyy9):

what will be the first step ?

jhonyy9 (jhonyy9):

what you know that l=2w

jhonyy9 (jhonyy9):

and the volum of rectangular container V=10 m^3

jhonyy9 (jhonyy9):

so do you know what is the formula of volum of a rectangular container ?

jhonyy9 (jhonyy9):

so if you dont like cooperating nothing i not can help to you sorry good luck bye

OpenStudy (anonymous):

i' sorry, i was on another problem, and posted this one to see response. the formula is V=BLW, but since 10 is the V, and L is 2b, it becomes V=(W)2(B^2). right?

jhonyy9 (jhonyy9):

what sign B there ?

OpenStudy (anonymous):

so B would be: \[\sqrt{5w}\] right? and then, it's asks for the costs, so would i then do: C=15(B)+9(S) C=15(B)+9(2L+2W) \[C=15(\sqrt{5w})+9(2(2\sqrt{5w})+2w)\] \[C=15\sqrt{5w}+36\sqrt{5w}+18w \] \[C=51\sqrt{5w}+18w\] then \[C \prime=\frac{ 51/2 }{ \sqrt{5w} } +10\] \[C \prime=0=\frac{ 25.5}{ \sqrt{5w} } +10\]

OpenStudy (anonymous):

what do you mean "sign B there"?

jhonyy9 (jhonyy9):

so there is V=BLW there what is B because you know that l is the lenght of base and w is the width of base so B what is ?

OpenStudy (anonymous):

ohhh ok. i though volume is ohhh... i did it wrong. it's suppose to be height isn't it.

jhonyy9 (jhonyy9):

so than for you can calcule this cost you need to know area of every sides ,yes ?

jhonyy9 (jhonyy9):

so than you need calcule the height ,yes ?